OpenStudy (anonymous):

A company sells boxes of cow bells (b) for $20 and boxes of air horns (h) for $30. They can make a batch of cow bells that fills 8 boxes or a batch of air horns that fills 9 boxes. The company only has 53 boxes. Which system of equations helps the company plan to make $275? A. b = 53 - h 8b + 9h = 53 B. 8b = 53 - 9h 20b + 30h = 275 C. b = 53 - h 20b + 30h = 275 D. 8b = 53 - 9h 20b - 30h = 275

OpenStudy (anonymous):

@texaschic101 @paki @ikram002p @DominicNg

OpenStudy (texaschic101):

cow bells (b) for 20 and air horns (h) for 30....plan to make 275 20b + 30h = 275 1 batch of airhorns ( 9 boxes) + 1 batch of cow bells (8 boxes) = 53 boxes 8b + 9h = 53 8b = 53 - 9h answer is B